Test scores.M any states have p rograms for assessing the skills of students in various grades. The Indiana Statewide Testing for Educational Progress (ISTEP) is one such p rogram.37In a recent y ear, 76,531, tenth-grade Indiana students took the English/language arts exam. The mean score was 572 and the standard deviation was 51. Assuming that these scores are ap p roximately Normally distributed, N(572, 51), use the 68–95–99.7 rule to give a range of scores that includes 95% of these students.
Answer #1

Using the 68 - 95 - 99.7 rule, the range of scores that will include 95% of the students = (μ-20. μ 2ơ) = (572-(2 * 51). 572 + (2 * 51))-(470.674) .